11. The interest rate on one-year risk-free bonds is 4.25 percent in the United States and 3.75 percent in Switzerland. The current exchange rate is $0.65 per Swiss frand. Suppose that you are a U.S, investor and you expect the Swiss franc to appreciate 2.75 percent over the next year. a. Calculate the foreign currency risk premium. b. Calculate the domestic currency return on the foreign bond, assuming that your cur- rency appreciation expectations are met
Answer #1

a) The interest rate differential =0.0425 −0.0375 =0.005, or 0.50%.

This implies that the Swiss franc trades at a forward premium of 0.50 percent. That is, the forward exchange rate is quoted at a premium of 0.50 percent over the spot exchange rate of $0.65 per Swiss franc.

The foreign currency risk premium =0.0275 −0.005 =0.0225, or 2.25%.

b) The domestic currency (U.S.$) return on the foreign bond is 6.5 percent. This can be calculated in one of two ways:

Domestic risk-free rate +Foreign currency risk premium =4.25% +2.25% =6.5%.

Foreign risk-free rate +Expected exchange rate movement =3.75% +2.75% =6.5%.